February 2012 Reports

To the main

29th February 2012 (Wednesday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • CERN : CASTOR problem to access some files : (GGUS:79629)
    • CERN : setting the variable TMPDIR : (GGUS:79685)

  • T1

28th February 2012 (Tuesday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • CERN : CASTOR problem to access some files : (GGUS:79629)
    • CERN : setting the variable TMPDIR : (GGUS:79685)

  • T1
    • PIC : Last patches installed at EMI WMS
    • PIC : Request for space token migration (GGUS:79305)
    • GridKa : Request for space token migration (GGUS:79303) "nearly finished"
    • SARA : Request for space token migration (GGUS:79307)

-

27th February 2012 (Monday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • CERN : CASTOR problem to access som efiles : (GGUS:79629)

  • T1
    • PIC : Last patches installed at EMI WMS
    • PIC : Request for space token migration (GGUS:79305)
    • GridKa : Request for space token migration (GGUS:79303) "nearly finished"
    • SARA : Request for space token migration (GGUS:79307)

  • T2
    • LAL (GGUS:79200).ongoing

24th February 2012 (Friday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • RAL : EMI WMS problem fixed

  • T1
    • PIC : Request for space token migration (GGUS:79305)
    • GridKa : Request for space token migration (GGUS:79303) "nearly finished"
    • SARA : Request for space token migration (GGUS:79307)

23rd February 2012 (Thursday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • CERN : EMI WMS problem fixed
2012-02-22 14:13:25 UTC WorkloadManagement/TaskQueueDirector/gLitePilotDirector INFO: Reference https://wms301.cern.ch:9000/GnF9HdDoA5-r5LwXWlT29Q for TaskQueue 2758122

2012-02-22 14:15:28 UTC WorkloadManagement/TaskQueueDirector/gLitePilotDirector INFO: Reference https://wms301.cern.ch:9000/Fpl3XHXTKY_-_ps3D-z9SQ for TaskQueue 2758135

Which means that problem is solved and LHCB is now successfully using this WMS.

The other WMS that need to updated to this patch are at least:

wms01.pic.es wms02.pic.es wms-lhcb.grid.cnaf.infn.it lcgwms01.gridpp.rl.ac.uk

22nd February 2012 (Wednesday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T1
    • PIC : Request for space token migration (GGUS:79305)
    • GridKa : Request for space token migration (GGUS:79303) "nearly finished"
    • SARA : Request for space token migration (GGUS:79307)

21st February 2012 (Tuesday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • SLS issue : someone remove some readonly volume yesterday around 3pm and one of our test use this volume was timing out . The problem is fixed now.

  • T1
    • IN2P3 : (GGUS:79356) : using the quue with more memory limit help to run our jobs.
    • PIC : Request for space token migration (GGUS:79305)
    • GridKa : Request for space token migration (GGUS:79303) "nearly finished"
    • SARA : Request for space token migration (GGUS:79307)

20th February 2012 (Monday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • NTR.

17th February 2012 (Friday)

Experiment activities:

User analysis, Validation productions for restripping

New GGUS (or RT) tickets

  • T0
    • NTR.

16th February 2012 (Thursday)

Experiment activities:

User analysis

New GGUS (or RT) tickets

  • T0
    • Problems transferring files to IN2P3. GGUS ticket (GGUS:79281) opened against IN2P3 - problem with network?

  • T1
    • IN2P3 : See above.
    • RAL : Continuing problems with publishing queue parameters / submission of jobs. GGUS ticket (GGUS:79283) submitted.

15th February 2012 (Wednesday)

Experiment activities:

User analysis

New GGUS (or RT) tickets

  • T0
    • lhcbDefault overload : Users informed and believe back to normal
    • Welcome move of srm-lhcb probe to LHCb-Disk

  • T1
    • RAL : Corrupted file. LHCb datamanagement informed. Also problems with batch server (internal ticket opened).

  • Other information

14th February 2012 (Tuesday)

Experiment activities:

User analysis

New GGUS (or RT) tickets

  • T1
    • RAL : Possible corrupted file. RAL Internal ticket opened.
    • GridKa : Slow staging of files

  • Other information
    • Proxy delegation bug on EMI WMS. LHCb would like to request that if possible not all WMS-es are upgraded until the bug is fixed

13th February 2012 (Monday)

Experiment activities:

User analysis

New GGUS (or RT) tickets

  • T1
    • IN2P3 : Jobs problem from last Friday - problem with LFC at IN2P3. Jobs recovered after the LFC there was rebooted. Currently a problem with ghost jobs at IN2P3 Cream CE (GGUS ticket 79164). Also a possibly related problem with LHCb pilots at IN2P3 not picking up jobs from the DIRAC task queue - currently analysing this problem.

  • Other information
    • Need to avoid having two Tier-1 sites simultaneously down if possible, at least for scheduled downtimes.
        • Note : 24 hours is minimum to declare a scheduled downtime.
    • Need CVMFS at GridKa (asap ...)
    • LHCb sees problems with proxy delegation with the EMI WMS at PIC

10th February 2012 (Friday)

Experiment activities:

MC11 Monte Carlo productions and user analysis

New GGUS (or RT) tickets

  • T1
    • IN2P3 : Jobs waiting for a long time after finishing - Tier-1 contact is investigating what could be happening there.

  • Other information
    • Need to avoid having two Tier-1 sites simultaneously down if possible, at least for scheduled downtimes.
    • Need CVMFS at GridKa (asap ...)
    • LHCb sees problems with proxy delegation with the EMI WMS at PIC

9th February 2012 (Thursday)

Experiment activities:

MC11 Monte Carlo productions and user analysis

New GGUS (or RT) tickets

  • T1
    • GRIDKA: Problems with LFC replication at GridKa - solved. Ticket closed.
    • RAL : Zombie jobs on CreamCE preventing direct job submission - these jobs have been killed. Question - how did they arise and can an automatic procedure be used to kill them?
    • IN2P3 : "srm authentication failed" trying to access some files. GGUS ticket 79074 opened and solved quickly. Waiting to see if it solves the problems seen by user jobs there.

8th February 2012 (Wednesday)

Experiment activities:

MC11 Monte Carlo productions and user analysis

New GGUS (or RT) tickets

  • T0
    • Looking forward to new hardware for DIRAC services.

  • T1
    • GRIDKA: Problems with LFC replication at GridKa. GGUS ticket 79014.
    • RAL : Zombie jobs on CreamCE preventing direct job submission which is LHCb preferred method of submission now. Submission via WMS-es working for now. GGUS ticket 78873.
    • IN2P3 : Problem with CVMFS when too many jobs start at the same time. Jobs hang setting up the environment (LbLogin / SetupProject) - this was before the last downtime.

7th February 2012 (Tuesday)

Experiment activities:

MC11 Monte Carlo productions and user analysis

New GGUS (or RT) tickets

  • T0
    • Problems contacting VOMS server from online cluster, most probably due to the firewall there.
    • Look forward to new hardware for DIRAC services.

  • T1
    • GRIDKA: Some backlog in merging jobs probably due to the problem with LFC (until it was banned yesterday). Upgrade of hardware of the LHCb Tier-1 VO-box at GridKa tomorrow.
    • RAL : Zombie jobs on CreamCE preventing direct job submission which is LHCb preferred method of submission now. Submission via WMS-es working for now. GGUS ticket 78873.
    • IN2P3 : Problem with CVMFS when too many jobs start at the same time. Jobs hang setting up the environment (LbLogin / SetupProject).

6s February 2012 (Monday)

Experiment activities:

MC11 Monte Carlo productions

New GGUS (or RT) tickets

  • T1
    • GRIDKA: Problem with LFC; changed to InActive in configuration

3d February 2012 (Friday)

Experiment activities:

MC11 Monte Carlo productions

New GGUS (or RT) tickets

  • T0
    • CERN Pilots aborted (GGUS:78893). Solved: LCG CEs removed from LHCb configuration, CREAMCEs fixed

2nd February 2012 (Thursday)

Experiment activities:

MC11 Monte Carlo productions

New GGUS (or RT) tickets

  • T0
  • T1

1st February 2012 (Wednesday)

Experiment activities:

MC11 Monte Carlo productions; Nothing to report

New GGUS (or RT) tickets

  • T0
  • T1

-- JoelClosier - 01-Apr-2012

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2012-04-01 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback